OpenStudy (anonymous):

if f(x)=x^(1/2)-3x and g(x)=2x^6-x^(1/2)-x find g(x) X f(x)

OpenStudy (anonymous):

@MelodicSymphony Is there a website i can plug this into?

OpenStudy (jdoe0001):

\(\bf f(x)=x^{\frac{1}{2}}-3x\qquad g(x)=2x^6-x^{\frac{1}{2}} \\ \quad \\ g(x)\times f(x)\implies (2x^6-x^{\frac{1}{2}})(x^{\frac{1}{2}}-3x) \)

OpenStudy (jdoe0001):

multiply away and recall your exponent rules http://www.math-play.com/image-exponents-rules.jpg

OpenStudy (anonymous):

would it be -6x^7+2x^(13/2) +3x^2+2x^(3/2)-x? @jdoe0001

OpenStudy (anonymous):

or all that but positive 6x^7

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

hmmm actaully my fraction additions are a bit off... should a bit more one sec

OpenStudy (jdoe0001):

\(\large { f(x)=x^{\frac{1}{2}}-3x\qquad g(x)=2x^6-x^{\frac{1}{2}} \\ \quad \\ g(x)\times f(x)\implies (2x^6-x^{\frac{1}{2}})(x^{\frac{1}{2}}-3x) \\ \quad \\ 2x^6x^{\frac{1}{2}}-2x^6\cdot 3x^1-x^{\frac{1}{2}}x^{\frac{1}{2}}+x^{\frac{1}{2}}\cdot 3x^1 \\ \quad \\ 2x^{6+\frac{1}{2}}-6x^{6+1}-x^{\frac{1}{2}+\frac{1}{2}}+3x^{\frac{1}{2}+1} \\ \quad \\ 2x^{\frac{13}{2}}-6x^7-x+3x^{\frac{3}{2}} }\) but yours is a bit close

OpenStudy (anonymous):

thank you!

OpenStudy (jdoe0001):

yw